A solenoid with 1000 turns/meter carries a current of 4 A. What is the resulting magnetic field?
Answer #1

n = 1000turn/m

I = 4A

B = u° nI = 4π×10-7 × 1000×4 = 16π×10^-4 T

=) B = 50.28×10^-4T = 5.03mT

Please give a thumb up. Thank you.